OpenStudy (anonymous):

Find the product of (4x − 3)(2x^2 − 7x + 1). A. 8x^3 − 22x^2 + 17x− 3 B, 8x^3 + 8x^2 + 4x − 3 C. 8x^3 − 34x^2 + 25x − 3 D. 8x^3 − 42x^2 + 25x − 3

OpenStudy (anonymous):

This is what i did (4x-3)(2x^2-7x+1) (4x-3)(2x^2) +(4x-3)(-7x)+(4x-3)(1) 8x^3-2x^2-28x^2+7x+4x-1 8x^3-30^4+11x^2-1

OpenStudy (ajprincess):

(4x-3)(2x^2)=8x^3-6x^2 u have to multiply 2x^2 by both 4x and -3

OpenStudy (ajprincess):

similarly (4x-3)(-7x)=(4x*-7x)-(3*-7x)=-28x^2-(-21x)=-28x^2+21x

OpenStudy (anonymous):

Sp what is the set up for this like when I'm combining like terms cause you just confused me complete haha

OpenStudy (anonymous):

@ajprincess